oxymethalone is an anabolic steroid used to treat anemia. an effective dose, which is administered over a 3 to 6 month period, is typically 2 mg/day for every kilogram of body mass. for a patient whose weight is 192.5 pounds, how many 50-mg tablets are needed for a daily dose of oxymethalone?
Answer #1

192.5lbs = 87.4kg

2mg/day/kg body weight = 2 x 87.4 mg/ day = 174.8mg/day

Everyday, number of 59mg tablets = 174.8/50 ~3 tablets

This is to be given for 3-6 months

Ie 90-180 days

= 15732- 31464mg total

This is to be given as 50mg tablets

So total number of tablets = 314.64-629.28 tablets for a period of 3-6mth